OpenStudy (anonymous):

the closed form of the sequence an: 2,7,12,17,... is ___. A. an=5n-3 B. an=2^n C. an=n+1 D. an=10n-8

OpenStudy (anonymous):

2,7,12,17,... This is an Arithmetic Sequence. The common difference is: +5

OpenStudy (anonymous):

2 + (5) = 7 7 + (5) = 12 12 + (5) = 17 17 + (5) = 22 and so on...

OpenStudy (anonymous):

I saw that but I plugged in some of the values into the equations but none of them worked... so I don't know if i'm doing it wrong or not

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

\[a _{n} = a _{1} + (n-1)d\] a1 is the first term. d is the common difference. \[a _{n} = 2 + (n-1)5\] \[a _{n} = 2 + 5n - 5\]

OpenStudy (anonymous):

\(a_1=2=5\times 1-3\)

OpenStudy (anonymous):

Can you finish it up from there @paigeRG

OpenStudy (anonymous):

and \(a_2=7=5\times 2-3\)

OpenStudy (anonymous):

oooooooooooooohhhhhhhhhh I get it now I was subtracting first instead of multiplying first thanks guys!!!

OpenStudy (anonymous):

yw :)